OpenStudy (anonymous):

Calculate the following integral \[\int\limits 1/(\sqrt (c^2 + m^2y^2)\]

OpenStudy (mimi_x3):

is it:\[\int \frac{1}{ \sqrt{c^2+ m^2y^2}} dy\]

OpenStudy (anonymous):

Yes it is

OpenStudy (mimi_x3):

substitute \[ y = \frac{ctan(u)}m\]

OpenStudy (mimi_x3):

\[ dy = \frac{cdusec^2(u)}{m}\]

OpenStudy (mimi_x3):

try it.

OpenStudy (anonymous):

So you get \[ \frac{ 1 }{(\sqrt{c^2 +m^2 *(\frac{ ctan(u) }{ m })^2}) } * \frac{ cdusec^2(u) }{ m }\]

OpenStudy (anonymous):

I don't really know how you can simplify that

OpenStudy (mimi_x3):

i think you can integrate \[\int sec(u)du\]

OpenStudy (anonymous):

Yes, thanks a lot

OpenStudy (mimi_x3):

:)